# A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v1

@@ 3.1 General case
Therefore, we cannot deduce a contradiction from the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map which is induced on the 00-the degree part of the homology from the DGA map as in Theorem 3.2.
@@
with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [20, Section 8].
We choose a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support such that the Lagrangian submanifold $P_{H}\coloneqq(\varphi^{1}_{H}\circ\psi)(\mathbb{R}^{3})$ intersects $L_{K_{0}}$ transversely. $P_{H}$ is oriented so that the diffeomorphism $\varphi^{1}_{H}\circ\psi\colon\mathbb{R}^{3}\to P_{H}$ preserves orientations. From the orientations of $P_{H}$ and $L_{K_{0}}$, a $\mathbb{Z}/2$-grading $|x|\in\mathbb{Z}/2$ is defined for every $x\in P_{H}\cap L_{K_{0}}$. See [19, Example 2.7, Section 2.d]. Note that the grading is shifted by $1$ if we use the opposite orientation of $P_{H}$. We also choose an almost complex structure $J$ such that $d\lambda_{\mathbb{R}^{3}}(\cdot,J\cdot)$ is a Riemannian metric on $T^{*}\mathbb{R}^{3}$ and $J(\partial_{p_{i}})=\partial_{q_{i}}$ for $i\in\{1,2,3\}$ outside a compact subset of $T^{*}\mathbb{R}^{3}$.
@@
Fix a base point $z_{0}\in L_{K_{0}}$. For any $x\in P_{H}\cap L_{K_{0}}$, we choose a path $c_{x}\colon[0,1]\to L_{K_{0}}$ such that $c_{x}(0)=z_{0}$ and $c_{x}(1)=x$. Let $c_{x}^{-1}$ denote the path $[0,1]\to L_{K_{0}}\colon t\mapsto c_{x}(1-t)$. For every $x,y\in P_{H}\cap L_{K_{0}}$ and a homology class $A\in H_{1}(L_{K_{0}})$, we define $\mathcal{M}_{(H,J),A}(x,y)$ to be the space of maps $u\colon\mathbb{R}\times[0,1]\to T^{*}\mathbb{R}^{3}$ satisfying:
  • su($\partial_{s}u(s,t)+J_{u(s,t)}(\partial_{t}u(s,t)-(X^{t}_{H})_{u(s,t)})=0$ for every $(s,t)\in\mathbb{R}\times[0,1]$.
@@
  • $A\in H_{1}(L_{K_{0}})$ is represented by a loop based at $z_{0}$ obtained by concatenating the three paths in $L_{K_{0}}$: $c_{x}$, the path $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto u(s,1)$ from $x$ to $y$, and $c_{y}^{-1}$.
@@
To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [19, Section 2.f]. We take the pair $(H,J)$ to be regular so that the moduli space $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is cut out transversely. The dimension of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is equal to $|x|-|y|-1$ modulo $2$. Moreover, by the Gromov compactness, the union of 00-dimensional components of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is a finite set. We define $n_{(H,J),A}(x,y)\in\mathbb{F}_{2}$ to be the modulo $2$ count of the number of the 00-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$.
@@
for every $y\in P_{H}\cap L_{K_{0}}$. By a standard argument using the Gromov compactification of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ for each $A\in H_{1}(L_{K_{0}})$ when its dimension is $1$, we can prove that $d_{J}\circ d_{J}=0$.
@@
Its isomorphism class as a $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-module is independent of the choice of the paths $\{c_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$. To see this claim, we choose another path $c^{\prime}_{x}$ instead of $c_{x}$ for every $x\in P_{H}\cap L_{K_{0}}$ and define $d^{\prime}_{J}$ by using $\{c^{\prime}_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$ in the same way as $d_{J}$. Take $m_{x}\in\mathbb{Z}$ such that $m_{x}[K_{0}]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating $c^{\prime}_{x}$ and $c_{x}^{-1}$. Then we can define an $\mathbb{F}_{2}$-linear isomorphism
@@
For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ in the statement of Proposition 2.7, let us set $\psi\coloneqq\varphi^{-1}$. Then, we can change the setup as follows: $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ have a clean intersection along $\psi(K_{1})$, where $K_{1}$ is an oriented knot in $\mathbb{R}^{3}$. We may assume that for sufficiently small $\varepsilon>0$, $\psi(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})\subset L_{K_{0}}$. We fix an orientation of $\psi(K_{1})$ via $\psi\colon K_{1}\to\psi(K_{1})$ and define $a\in\mathbb{Z}$ by $[\psi(K_{1})]=a[K_{0}]\in H_{1}(L_{K_{0}})$. Let $(-1)^{\sigma}$ denote the sign of orientation of the open embedding $\psi\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$. Then, the claim of the proposition is equivalent to that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$.
@@
Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [18, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $||f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold:
@@
since $[\psi(K_{1})]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating the two paths: $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{1}(-s)$ and $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{2}(s)$. From this observation, the differential $d_{J_{g}}\colon\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{1}\to\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{0}$ is given by
@@
Here, $\sigma_{\psi}\in\mathbb{Z}/2$ is the shift of degrees. Note that $\sigma_{\psi}$ depends only on the orientation of $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$ induced by the open embedding $\psi\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$ and the fixed orientation of $L_{K_{0}}$. As $\mathbb{F}_{2}$-vector spaces,
@@
From the definition of the orientation sign $(-1)^{\sigma}$, the shift of degrees $\sigma_{\psi}$ is equal to $\sigma_{\operatorname{id}}+\sigma$ modulo $2$. From (15), (17) and Theorem A.1, it follows that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$. ∎

### v2

@@ 3.2.3 Application of $\mathbb{Q}$-valued augmentations
Therefore, we cannot deduce a contradiction from the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map between the 0-th homology which is induced from the DGA map as in Theorem 3.2.
3.3 Proof of Theorem 1.1 for $K^{\prime}\#4_{1}$
Fix any $B\in B_{n}$ whose closure gives a knot $K$. We consider a braid on $(n+3)$ strands
$B^{\#}\coloneqq\sigma_{n}\left(\sigma_{n+1}(\sigma_{n+2})^{-1}\sigma_{n+1}(\sigma_{n+2})^{-1}\right)B\in B_{n+3}.$
Then, the closure of $B^{\#}$ is the connected sum of $K$ and the figure-eight knot $4_{1}$.
Proposition 3.17.
We define $P\in\mathbb{Z}[\mu^{\pm},U^{\pm}][T]$ by
$\begin{split}P(T)\coloneqq&-(1-\mu)(U-\mu)+\left(-2+2\mu+\mu^{2}+(\mu^{-1}-1-\mu)U-\mu^{3}U^{-1}\right)T\\ &+(-2\mu^{2}+\mu U)T^{2}-\mu UT^{3}.\end{split}$ (25)
Let $\mathbb{k}$ be a field. Then, for any augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#4_{1})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)$, the polynomial $\left.P\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
Proof.
First, let us compute $\phi_{B^{\#}}(a_{n+2,*})$ and $\phi_{B^{\#}}(a_{n+3,*})$. Since $\phi_{\sigma_{k}}(a_{n+2,*})=a_{n+2,*}$ and $\phi_{\sigma_{k}}(a_{n+3,*})=a_{n+3,*}$ for every $k\in\{1,\dots,n-1\}$, we have $\phi_{B}(a_{n+2,*})=a_{n+2,*}$ and $\phi_{B}(a_{n+3,*})=a_{n+3,*}$. Referring to (10), we compute that
$(\phi_{\sigma_{n+2}})^{-1}\colon\begin{cases}a_{n+1,*}\mapsto a_{n+1,*},\\ a_{n+2,*}\mapsto a_{n+3,*},\\ a_{n+3,*}\mapsto a_{n+2,*}-a_{n+2,n+3}a_{n+3,*},\\ a_{n+1,n+3}\mapsto a_{n+1,n+2}-a_{n+1,n+3}a_{n+3,n+2},\end{cases}$
and
${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\phi_{\sigma_{n+2}})^{-1}}$${a_{n+3,*}}$${a_{n+3,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\phi_{\sigma_{n+2}})^{-1}}$${a_{n+2,*}-a_{n+2,n+3}a_{n+3,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+3,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}-a_{n+1,n+3}a_{n+3,*}}$$\scriptstyle{(\phi_{\sigma_{n+2}})^{-1}}$${a_{n+2,*}-a_{n+2,n+3}a_{n+3,*}}$$\scriptstyle{(\phi_{\sigma_{n+2}})^{-1}}$${a_{n+1,*}-M^{\prime\prime}(a_{n+2,*}-a_{n+2,n+3}a_{n+3,*})}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}-a_{n+1,n+3}a_{n+3,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${\begin{array}[]{l}a_{n+2,*}-a_{n+2,n+1}a_{n+1,*}\\ -M^{\prime}(a_{n+1,*}-a_{n+1,n+3}a_{n+3,*})\end{array}}$$\scriptstyle{\phi_{\sigma_{n}}}$${a_{n,*}-a_{n,n+3}a_{n+3,*},}$$\scriptstyle{\phi_{\sigma_{n}}}$${a_{n+2,*}-a_{n+2,n}a_{n,*}-M(a_{n,*}-a_{n,n+3}a_{n+3,*}),}$
where
$M^{\prime\prime}$ $\coloneqq(\phi_{\sigma_{n+2}})^{-1}(a_{n+1,n+3})=a_{n+1,n+2}-a_{n+1,n+3}a_{n+3,n+2},$
$M^{\prime}$ $\coloneqq\phi_{\sigma_{n+1}}(M^{\prime\prime})=-a_{n+2,n+1}-(a_{n+2,n+3}-a_{n+2,n+1}a_{n+1,n+3})a_{n+3,n+1},$
$M$ $\coloneqq\phi_{\sigma_{n}}(M^{\prime})=-a_{n+2,n}-(a_{n+2,n+3}-a_{n+2,n}a_{n,n+3})a_{n+3,n}.$
As a consequence, we obtain
$\phi_{B^{\#}}(a_{n+2,*})$ $=a_{n,*}-a_{n,n+3}a_{n+3,*},$
$\phi_{B^{\#}}(a_{n+3,*})$ $=(-a_{n+2,n}-M)a_{n,*}+a_{n+2,*}+(Ma_{n,n+3})a_{n+3,*},$
for
$M=-a_{n+2,n}(1-a_{n,n+3}a_{n+3,n})-a_{n+2,n+3}a_{n+3,n}.$ (26)
The $(n+2)$-th and $(n+3)$-th row of $\bm{\Phi}^{L}_{B^{\#}}$ are given by
$\begin{pmatrix}0&\cdots&0&1&0&0&-a_{n,n+3}\\ 0&\cdots&0&-a_{n+2,n}-M&0&1&Ma_{n,n+3}\end{pmatrix}$.
Let $c_{i,j}$ denote the $(i,j)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$. Then,
$c_{n+2,n}$ $=\widehat{\mathbb{A}}_{n+2,n}-(1\cdot\mathbb{A}_{n,n}-a_{n,n+3}\mathbb{A}_{n+3,n})$
$=-\mu a_{n+2,n}-(1-\mu)-\mu a_{n,n+3}a_{n+3,n},$
$c_{n+2,n+3}$ $=\widehat{\mathbb{A}}_{n+2,n+3}-(1\cdot\mathbb{A}_{n,n+3}-a_{n,n+3}\mathbb{A}_{n+3,n+3})$
$=Ua_{n+2,n+3}-a_{n,n+3}+(1-\mu)a_{n,n+3}$
$=Ua_{n+2,n+3}-\mu a_{n,n+3},$
$c_{n+3,n}$ $=\widehat{\mathbb{A}}_{n+3,n}-((-a_{n+2,n}-M)\mathbb{A}_{n,n}+1\cdot\mathbb{A}_{n+2,n}+Ma_{n,n+3}\mathbb{A}_{n+3,n})$
$=-\mu a_{n+3,n}+(1-\mu)(a_{n+2,n}+M)+\mu a_{n+2,n}+\mu Ma_{n,n+3}a_{n+3,n}$
$=-\mu a_{n+3,n}+a_{n+2,n}+M(1-\mu+\mu a_{n,n+3}a_{n+3,n}),$
$c_{n+3,n+3}$ $=\widehat{\mathbb{A}}_{n+3,n+3}-((-a_{n+2,n}-M)\mathbb{A}_{n,n+3}+1\cdot\mathbb{A}_{n+2,n+3}+Ma_{n,n+3}\mathbb{A}_{n+3,n+3})$
$=(U-\mu)+(a_{n+2,n}+M)a_{n,n+3}-a_{n+2,n+3}-(1-\mu)Ma_{n,n+3}$
$=(U-\mu)+a_{n+2,n}a_{n,n+3}-a_{n+2,n+3}+\mu Ma_{n,n+3}.$
Let us denote $X=a_{n,n+3}$ and $Y=a_{n+3,n}$. The ideal $\mathcal{I}_{B^{\#}}$ contains $\pi_{n+3}(c_{i,j})$ for every $i,j\in\{1,\dots,n+3\}$, where $\pi_{n+3}$ is defined as (13). In particular, $\mathcal{I}_{B^{\#}}$ contains
$\mu^{-1}\cdot\pi_{n+3}(c_{n+2,n})$ $=-a_{n+2,n}-\mu^{-1}+1-XY,$
$U^{-1}\cdot\pi_{n+3}(c_{n+2,n+3})$ $=a_{n+2,n+3}-\mu U^{-1}X.$
Replacing $a_{n+2,n}$ with $\mu^{-1}+1-XY$ and $a_{n+2,n+3}$ with $\mu U^{-1}X$, we obtain from $\pi_{n+3}(M)$
$\bar{M}$ $\coloneqq-(-\mu^{-1}+1-XY)(1-XY)-\mu U^{-1}XY$
$=\mu^{-1}(1-\mu)+(\mu^{-1}-\mu U^{-1})XY-(XY)^{2},$
and by the same replacement, we obtain from $\pi_{n+3}(c_{n+3,n})$ and $\pi_{n+3}(c_{n+3,n+3})$
$F$ $\coloneqq-\mu Y+(-\mu^{-1}+1-XY)+\bar{M}(1-\mu+\mu XY),$
$G$ $\coloneqq(U-\mu)+(-\mu^{-1}+1-XY)X-\mu U^{-1}X+\mu\bar{M}X$
$=(U-\mu)+(-\mu^{-1}+1-\mu U^{-1}-XY)X+\mu\bar{M}X.$
Then, both $F$ and $G$ are contained in $\mathcal{I}_{B^{\#}}$.
We focus on the first terms $\mu Y$ and $(U-\mu)$ of $F$ and $G$. Then, we compute
$(U-\mu)\cdot F+\mu Y\cdot G=$ $(U-\mu)(-\mu^{-1}+1-XY)+(U-\mu)\bar{M}(1-\mu+\mu XY)$
$+\mu(-\mu^{-1}+1-\mu U^{-1}-XY)XY+\mu^{2}\bar{M}XY$
$=$ $-\mu^{-1}(1-\mu)(U-\mu)+(-1+2\mu-U-\mu^{2}U^{-1})XY-\mu(XY)^{2}$
$+\bar{M}\left((1-\mu)(U-\mu)+\mu UXY\right)$
and
$\bar{M}\left((1-\mu)(U-\mu)+\mu UXY\right)$
$=$ $\left(\mu^{-1}(1-\mu)+(\mu^{-1}-\mu U^{-1})XY-(XY)^{2}\right)\left((1-\mu)(U-\mu)+\mu UXY\right)$
$=$ $\mu^{-1}(1-\mu)^{2}(U-\mu)+(1-\mu)(U\mu^{-1}-1-\mu+U+\mu^{2}U^{-1})XY$
$+(\mu-2\mu^{2}+\mu U)(XY)^{2}-\mu U(XY)^{3}.$
Therefore,
$(U-\mu)\cdot F+\mu Y\cdot G$
$=$ $-(1-\mu)(U-\mu)+\left(-2+2\mu+\mu^{2}+(\mu^{-1}-1-\mu)U-\mu^{3}U^{-1}\right)XY$
$+(-2\mu^{2}+\mu U)(XY)^{2}-\mu U(XY)^{3}.$
If we define $P\in\mathbb{Z}[\mu^{\pm},U][T]$ by (25), then
$(U-\mu)\cdot F+\mu Y\cdot G=P(XY),$
which implies that $P(XY)$ is contained in $\mathcal{I}_{B^{\#}}$. Therefore, we have a well-defined unital $R$-algebra map
$J\colon R[T]/(P)\to R[a_{i,j}\mid 1\leq i,j\leq n+3,\ i\neq j]/\mathcal{I}_{B^{\#}}=HC^{\mathrm{ab}}_{0}(K\#4_{1})$
which maps $T$ to $XY=a_{n,n+3}a_{n+3,n}$. Now, the proposition follows from Lemma 3.4. ∎
Remark 3.18.
By the substitution $U=1$,
$\left.P\right|_{U=1}(T)=$ $-(1-\mu)^{2}+(\mu^{-1}-3+\mu+\mu^{2}-\mu^{3})T+(-2\mu^{2}+\mu)T^{2}-\mu T^{3}$
$=$ $-(T+\mu^{-1}(1-\mu)^{2})\left(\mu T^{2}+(-1+\mu+\mu^{2})T+\mu\right).$
Therefore, $\left.P\right|_{U=1}\in\mathbb{Z}[\mu^{\pm}][T]$ has a root $T=-\mu^{-1}(1-\mu)^{2}$. Hence, Proposition 3.17 is trivial if an augmentation $\varepsilon$ satisfies $\varepsilon(U)=1$.
Theorem 3.19.
Let $K_{0}=K\#4_{1}$, where $K$ is an arbitrary knot in $\mathbb{R}^{3}$. Then, there is no $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along the unknot in $\mathbb{R}^{3}$.
Proof.
For a knot $K_{0}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$. Assuming that $K_{0}=K\#4_{1}$, we deduce a contradiction as follows: By Proposition 3.5, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{Q}$ such that $\varepsilon(\mu)=-1$ and $\varepsilon(U)=-2$. Then, it follows from Proposition 3.17 that the polynomial $\bar{P}\coloneqq\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$. On the other hand, from the definition of $P$,
$\bar{P}(T)=2-\frac{3}{2}T-2T^{3}$
Multiplying by $2$, we obtain a polynomial with integer coefficients
$4-3T-4T^{3}.$
By the rational root theorem, the only possible rational roots of this polynomial are of the form $\frac{p}{q}$, where $p\in\mathbb{Z}$ is a divisor of $4$ and $q\in\mathbb{Z}$ is a divisor of 4-4. The set of such rational numbers consists of the integers
$\pm 1,\pm 2,\pm 4,$
and those which are not integers
$\pm\frac{1}{2},\pm\frac{1}{4}.$
We note that $2\cdot\bar{P}(T)=4-T(4T^{2}+3)$. For every $n\in\{\pm 1,\pm 2,\pm 4\}$, the integer $4n^{2}+3$ is not a divisor of $4$, which implies that $2\cdot\bar{P}(n)\neq 0$. One can also check that
$\begin{array}[]{llll}\bar{P}(\frac{1}{2})=1,&\bar{P}(-\frac{1}{2})=3,&\bar{P}(\frac{1}{4})=\frac{51}{32},&\bar{P}(-\frac{1}{4})=\frac{77}{32}.\end{array}$
Therefore, we can conclude that $\bar{P}$ does not have a root in $\mathbb{Q}$, which contradicts Proposition 3.5 and Proposition 3.17. ∎
@@
with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [19, Section 8].
We choose a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support such that the Lagrangian submanifold $P_{H}\coloneqq(\varphi^{1}_{H}\circ\psi)(\mathbb{R}^{3})$ intersects $L_{K_{0}}$ transversely. $P_{H}$ is oriented so that the diffeomorphism $\varphi^{1}_{H}\circ\psi\colon\mathbb{R}^{3}\to P_{H}$ preserves orientations. From the orientations of $P_{H}$ and $L_{K_{0}}$, a $\mathbb{Z}/2$-grading $|x|\in\mathbb{Z}/2$ is defined for every $x\in P_{H}\cap L_{K_{0}}$. See [18, Example 2.7, Section 2.d]. Note that the grading is shifted by $1$ if we use the opposite orientation of $P_{H}$. We also choose an almost complex structure $J$ such that $d\lambda_{\mathbb{R}^{3}}(\cdot,J\cdot)$ is a Riemannian metric on $T^{*}\mathbb{R}^{3}$ and $J(\partial_{p_{i}})=\partial_{q_{i}}$ for $i\in\{1,2,3\}$ outside a compact subset of $T^{*}\mathbb{R}^{3}$.
@@
Fix a base point $z_{0}\in L_{K_{0}}$. For any $x\in P_{H}\cap L_{K_{0}}$, we choose a path $c_{x}\colon[0,1]\to L_{K_{0}}$ such that $c_{x}(0)=z_{0}$ and $c_{x}(1)=x$. Let $c_{x}^{-1}$ denote the path $0,1]\to L_{K_{0}}\colon t\mapsto c_{x}(1-t)$. For every $x,y\in P_{H}\cap L_{K_{0}}$ and a homology class $A\in H_{1}(L_{K_{0}})$, we define $\mathcal{M}_{(H,J),A}(x,y)$ to be the space of maps $u\colon\mathbb{R}\times[0,1]\to T^{*}\mathbb{R}^{3}$ satisfying:
  • $\partial_{s}u(s,t)+J_{u(s,t)}(\partial_{t}u(s,t)-(X^{t}_{H})_{u(s,t)})=0$ for every $(s,t)\in\mathbb{R}\times[0,1]$.
@@
  • $A\in H_{1}(L_{K_{0}})$ is represented by a loop based at $z_{0}$ obtained by concatenating the three paths in $L_{K_{0}}$: $c_{x}$, the path $\infty,\infty]\to L_{K_{0}}\colon s\mapsto u(s,1)$ from $x$ to $y$, and $c_{y}^{-1}$.
@@
To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [18, Section 2.f]. We take the pair $(H,J)$ to be regular so that the moduli space $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is cut out transversely for every $x,y\in P_{H}\cap L_{K_{0}}$ such that $x\neq y$. The dimension of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is equal to |x||y|1|x|-|y|-1 modulo $2$. Moreover, by the Gromov compactness, the union of 0-dimensional components of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is a finite set. We define $n_{(H,J),A}(x,y)\in\mathbb{F}_{2}$ to be the modulo $2$ count of the number of the 0-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$.
@@
for every $y\in P_{H}\cap L_{K_{0}}$. By a standard argument using the Gromov compactification of the $1$-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ for each $A\in H_{1}(L_{K_{0}})$, we can prove that $d_{J}\circ d_{J}=0$.
@@
We claim that its isomorphism class as a $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-module is independent of the choice of the paths $\{c_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$. To see this, we choose another path $c^{\prime}_{x}$ instead of $c_{x}$ for every $x\in P_{H}\cap L_{K_{0}}$ and define $d^{\prime}_{J}$ by using $\{c^{\prime}_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$ in the same way as $d_{J}$. Take $m_{x}\in\mathbb{Z}$ such that $m_{x}[K_{0}]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating $c^{\prime}_{x}$ and $c_{x}^{-1}$. Then we can define an $\mathbb{F}_{2}$-linear isomorphism
@@
For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ in the statement of Proposition 2.7, let us set $\psi\coloneqq\varphi^{-1}$. Then, we can change the setup as follows: $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ have a clean intersection along $\psi(K_{1})$, where $K_{1}$ is an oriented knot in $\mathbb{R}^{3}$. We may assume that for sufficiently small $\varepsilon>0$, $\psi(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})\subset L_{K_{0}}$. We fix an orientation of $\psi(K_{1})$ via the diffeomorphism $\left.\psi\right|_{K_{1}}\colon K_{1}\to\psi(K_{1})$ and define $a\in\mathbb{Z}$ by $\psi(K_{1})]=a[K_{0}]\in H_{1}(L_{K_{0}})$. Let $(-1)^{\sigma}$ denote the sign of orientation of the embedding $\left.\psi\right|_{L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}}\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$. Then, the claim of the proposition is equivalent to that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$. Since $\sigma\equiv 0\mod 2$ is already shown in [16, Lemma 4.4], we only give a proof for $a\in\{1,-1\}$ (see also Remark A.2 below).
@@
Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [17, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold:
@@
since $\psi(K_{1})]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating the two paths: $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{1}(-s)$ and $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{2}(s)$. From this observation, the differential $d_{J_{g}}\colon\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{1}\to\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{0}$ is given by
@@
Here, $\sigma_{\psi}\in\mathbb{Z}/2$ is the shift of degrees. As $\mathbb{F}_{2}$-vector spaces,
@@
From (29), (31) and Theorem A.1, it follows that $\sigma_{\psi}\equiv\sigma_{\operatorname{id}}\mod 2$ and $a\in\{1,-1\}$. ∎
Remark A.2.
The fact $\sigma\equiv 0\mod 2$ was proved in [16, Lemma 4.4] by introducing $\mathbb{Z}$-gradings of Lagrangian submanifolds which are lifts of orientations of Lagrangian submanifolds. Analogous to [16, Remark 2.9, Lemma 4.4], by an argument on $\mathbb{Z}/2$-grading in the present case, we can deduce $\sigma\equiv 0\mod 2$ from $\sigma_{\psi}\equiv\sigma_{\operatorname{id}}\mod 2$.